at an electronics store they sell about 730 cassette players a year and about 180 CD players a year. If the rate of cassette player sales decreases by 50each year and the rate of CD player sales increases by 60 each year, in how many years will the CD player sales be twice as much as the cassette player sales? use a table to determine you answer.
Answer by stanbon(75887)

at an electronics store they sell about 730 cassette players a year
and about 180 CD players a year.
If the rate of cassette player sales decreases by 50 each year and the rate of CD player sales increases by 60 each year,
----------------------
Cassette Equation: salesC(x)= 730 -50x, where x is the number of years from now.
----------------------
CD Equation: salesCD(x)= 180 + 60x, where x is the number of years from now.
----------------------
In how many years will the CD player sales be twice as much as the cassette player sales?
180 + 60x = 2(730-50x)
180 + 60x = 1460 - 100x
160x = 1280
x = 8 years
=================
Cheers,
Stan H.
use a table to determine you answer.